# Sonja Hohloch  &\&&  Pedro Santos

### v1

@@ 3.1. Generalization of the method of probes
Let $F:(M,\omega)\rightarrow B$ be a toric fibration. Suppose that there exists an affine embedding of $([0,R]\times[0,h],\mathcal{A}_{0})$ into $(B,\mathcal{A})$ where $\mathcal{A}$ is the integral affine structure induced by $F$ on $B$, and $R>0$ and $0<h<\frac{R}{2}$. Then the set $F^{-1}([0,h]\times[0,h])\subset M$ is displaceable.
@@
Let $F:(M,\omega)\rightarrow B$ be a toric fibration. Suppose that there exists an affine embedding of $([0,R]\times[0,a],\mathcal{A}_{0})$ into $(B,\mathcal{A})$ where $\mathcal{A}$ is the integral affine structure induced by $F$ on $B$, and $R,a>0$, $0<h<\frac{R}{2}$. Then the set $F^{-1}([0,h]\times[0,a])\subset M$ is displaceable.
@@ 3.2. Nondisplaceability of focus-focus fibers with higher multiplicity
Let $c$ be a focus-focus value of a semitoric system $F$ on a $4$-dimensional symplectic manifold $(M,\omega)$, and let $k$ be the number of focus-focus point on the fiber $F^{-1}(c)$, often called the multiplicity of the focus-focus fiber $F^{-1}(c)$. Intuitively, a fiber with $k$ focus-focus points is often seen as a torus with $k$ pinches. Alternatively the fiber $F^{-1}(c)$ can be seen as a closed chain of $k$ spheres joint at the poles, see Zung [zungarnold] and Vũ Ngǫc [vu2000bohr, Section $6$]. In particular, these are Lagrangian spheres. If $k=1$ the sphere is immersed, and if $k>1$ the spheres are embedded.
@@
Recall the notion of intersection number and Euler characteristic (for a reference see Hirsch [hirsch2012differential, Section $5.2$]).
Consider an embedded Lagrangian sphere $i:L\rightarrow(M,\omega)$ where $(M,\omega)$ is a $4$-dimensional symplectic manifold. Using Weinstein’s Lagrangian neighborhood theorem, we identify a neighborhood of $i(L)$ with the cotangent bundle of $S^{2}$. Therefore the intersection number $L\cdot L$ is such that $|L\cdot L|=\chi(S^{2})\neq 0$. Now we prove that $i(L)$ is not topologically displaceable by a smooth map $g:L\rightarrow M$ homotopic to the inclusion $i$. Let us argue by contradiction. Assume that $g(i(L))\cap i(L)=\emptyset$ and $g\simeq i$. In particular, $g$ and $i$ are transverse, and since $g(i(L))\cap i(L)=\emptyset$ we have $L\cdot L=\#(g,L)=0$, where #$\#(g,L)$ is the intersection number of the map $g$ with $L$. Therefore we obtain a contradiction. Since $L$ is not topologically displaceable, in particular it is not displaceable by a Hamiltonian diffeomorphism. ∎
@@
Therefore studying (non)displaceability of focus-focus fibers is reduced to the case $k=1$, which is the generic case, see Smirnov [smirnov2013focus]. There are also explicit examples for higher multiplicity. For the case $k=2$ see for instance De Meulenaere & Hohloch [de2021family] and Section 4.3.
@@ 3.3. Displacing focus-focus fibers
Now assume that the corner introduced by the focus-focus fiber is a Delzant corner, and let $A$ be the integral affine transformation that takes it into the standard Delzant corner. By Remark 2.20 we can apply a nodal trade (which is defined in Lemma 2.15). Recall the affine distance from Equation (2.1) and denote the affine length by aff. Without loss of generality assume that $\text{aff}(v_{1})\geq\text{aff}(v_{2})$. Suppose that $h<\text{aff}(v_{2})$, and that there exists $R>0$ such that $R<\text{aff}(v_{1})$ and $0<2h<R$. Furthermore, suppose that the rectangle $[0,R]\times[0,h]$ starting at $(0,0)$ sits inside $A(\Delta_{\vec{\epsilon}})$, as is sketched in Figure 3.2. Then we obtain the following result:
@@
Figure 3.2. Rectangle $[0,R]\times[0,h]$ sitting inside $A(\Delta_{\vec{\epsilon}})$. The red dot represents the focus-focus value.
@@
In this section we consider (non)displaceability of fibers of the coupled spin oscillator, Subsection 4.1, the coupled angular momenta, Subsection 4.2, and a semitoric system on the symplectic manifold given by the octagon, Subsection 4.3.
@@
Let $\rho_{1},\rho_{2}>0$ be positive constants. Consider the product manifold $M=S^{2}\times\mathbb{R}^{2}$ with symplectic form $\omega=\rho_{1}\omega_{S^{2}}\bigoplus\rho_{2}\omega_{\mathbb{R}^{2}}$, where $\omega_{S^{2}}$ and $\omega_{\mathbb{R}^{2}}$ are the standard symplectic structures on the unit sphere and the Euclidean space respectively. Let $(x,y,z)$ be Cartesian coordinates on the unit sphere $S^{2}\subset\mathbb{R}^{3}$ and $(u,v)$ Cartesian coordinates on the plane $\mathbb{R}^{2}$.
@@
Coupled spin-oscillators are semitoric systems, for more details and background see Pelayo & Vũ Ngǫc [pelayo2012hamiltonian]. The system has exactly one focus-focus singularity at the point $m:=(0,0,1,0,0)$. In order to understand the displaceability properties of the fibers of these systems we look at a representative of the polytope invariant, see Figure 4.1. These were computed in Pelayo & Vũ Ngǫc  [pelayo2012hamiltonian] and Alonso [alonso2019symplecticthesis, Section $5.1$].
Figure 4.1. Representative of the polytope invariant of the coupled-spin oscillator with $\vec{\epsilon}=+1$. The red dot represents the focus-focus value and is placed at $(0,0)$. Notice that the polytope is unbounded.
@@
Let $R_{1},R_{2}\in\mathbb{R}^{>0}$ and $t\in[0,1]$ a parameter. Consider the product manifold $M:=S^{2}\times S^{2}$ with symplectic form $\omega:=-(R_{1}\omega_{S^{2}}\bigoplus R_{2}\omega_{S^{2}})$ where $\omega_{S^{2}}$ is the standard symplectic form on $S^{2}$. Let $(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2})$ be Cartesian coordinates on $M$ induced from the ambient ${\mathbb{R}}^{3}\times{\mathbb{R}}^{3}$. The coupled angular momenta system is a family of $4$-dimensional completely integrable systems $(M,\omega,F_{t}:=(L,H_{t}))$, where the smooth functions $L,H_{t}:M\rightarrow\mathbb{R}$ are given by
@@
The system has four fixed points located at $(0,0,\pm 1,0,0,\pm 1)$. Three are of elliptic-elliptic type, namely $(0,0,\pm 1,0,0,1)$, but $(0,0,1,0,0,-1)$ is of focus-focus type for $t\in\ ]t^{-},t^{+}[$ and degenerate for $t\in\{t^{-},t{+}\}$, where
@@
Note that at the parameter $t=\frac{1}{2}$ there is always a focus-focus singularity. For more details and proofs of these results, see for instance Alonso [alonso2019symplecticthesis, Section 6] and Alonso & Dullin & Hohloch [alonso2019symplectic].
@@
Note that the map $f$, defined as the composition of a horizontal translation with the straightening homeomorphism $f_{\vec{\epsilon}}$ of a representative of the polytope invariant, maps the vertical line $x=0$ in $F_{t}(M)$ to the vertical line $x=R_{2}-R_{1}$ in $f(F_{t}(M))$, see Alonso & Dullin & Hohloch [alonso2019symplectic].
@@
Moreover, for $b\in\mathbb{R}^{>0}$, we have $[-b-2t,-b]\cap\{b\}=\emptyset$, hence the desired fibers are displaceable by $\Psi$. Note that the focus-focus value is $(0,1-2t)\in{\mathbb{R}}^{2}$. Hence, for $t<\frac{1}{2}$, the focus-focus fiber $F_{t}^{-1}(0,1-2t)$ is displaceable. ∎
Let us now inquire about the (non)displaceability of the focus-focus fiber for $t\geq\frac{1}{2}$. First we need to establish some facts about the height invariant. From now on $h:=h_{t}$ is the height invariant associated with the Kepler problem $(M,\omega,F_{t})$. In Alonso & Dullin & Hohloch [alonso2019symplectic] the authors compute the height invariant of the Kepler problem and show that it is given by
@@
Figure 4.3. Plot of $\tilde{h}$ for $t\in\ ]t^{-},t^{+}[$ in blue. Plot of the constant line equal $1$ in yellow. Plot of the constant line equal 00 in green.
@@
Step 1: Displaceability on the segment of the eigenline on the focus-focus value. Due to the choice of $t_{0}$ we have that for $t^{-}<t<t_{0}$, $\tilde{h}(t)>1$ and thus $2R-h<R$. Hence we can apply Lemma 3.5 in Figure 4.5 to conclude that the focus-focus fiber and every fiber of the form $F_{t}^{-1}(0,b)$, where $(0,b)\in F_{t}(M)$ with $b>1-2t$, is displaceable.
@@
We are left with understanding the cases $t\geq t_{0}$. In Theorem 4.12 we will obtain a complete description of the (non)displaceability properties of the Kepler problem for $t=t_{0}$: there exists only one nondisplaceable fiber, and it is given by the focus-focus fiber, in particular it is a stem and hence nondisplaceable by symplectomorphisms. For $t>t_{0}$ we will see that there exist an infinite amount of nondisplaceable fibers, noticing that these fibers are equivalent to fibers already studied in the literature, see Auroux [auroux2007mirror]. In particular the focus-focus fiber is nondisplaceable. This is done in Section 4.2.3.
@@
In order to obtain further displaceability results we need to look at another representative of the polytope invariant for the Kepler problem. We use the one given by Figure 4.6, for more details see Alonso & Dullin & Hohloch [alonso2019symplectic].
@@
Let $(0,y)\in f_{\vec{\epsilon}}(F_{t}(M))$ and $y>\frac{4}{3}R-h$. Notice that the vector $v:=(1,0)$ is integrally transverse to the facet given by $y=x+(2R-h)$. Therefore we can construct the probe $(x+k,y)$ where $y=x+(2R-h)>0$. The affine distance of this probe is $2R-\frac{(y+h)}{2}$ since it will end at the facet given by $y=2x-h$. Due to Lemma 2.22 the $(0,y)$ such that $(0,y)\in f_{\vec{\epsilon}}(F_{t}(M))$ and $y>\frac{4}{3}R-h$ give displaceable fibers, since they lie less than halfway along the probe. ∎
@@
Consider the Kepler problem for $t_{0}\leq t<1$. Let $f_{\vec{\epsilon}}$ be the straightening homeomorphism associated with the polytope invariant given by Figure 4.6. Then the fibers $F_{t}^{-1}(f_{\vec{\epsilon}}^{-1}(0,y))$ with $(0,y)\in f_{\vec{\epsilon}}(F_{t}(M))$ and $y>R-h$ of the Kepler problem are displaceable.
@@
The fibers we consider here correspond to fibers that sit above the segment of the eigenline $(x,x)$ in Figure 4.5 with $x<2R-h$. In the literature these are known as Chekanov-type tori, see Auroux [auroux2007mirror].
By applying a nodal slide, see Proposition 2.14, we can make the segment of the eigenline of Figure 4.6 as small as possible while preserving the fiber we are considering. This corresponds to considering the fiber $(x,x)$ in the polytope invariant, given by Figure 4.6, of the system for a $\tilde{t}>t$. Therefore the fiber over $(x,x)$ is displaceable by the probe that starts in the facet $y=x+(2R-h)$ with direction $(0,-1)$. Hence we obtain our desired result. ∎
@@
Since the fiber for the Kepler problem $(M,\omega,F_{t_{0}})$ is actually monotone, one could use the methods of Hong & Kim & Lau [hong2023immersed] or Rizzel & Ekholm & Tonkonog [rizell2018refined] to obtain the potential function for the focus-focus fiber, and then obtain a nondisplaceability result by computing the critical points of the potential function. However the result proven in Theorem 4.12 is stronger since we prove that the focus-focus fiber is a stem.
@@
Let $N\in\mathbb{N}$ and $R>0$. Then, for $11-\frac{1}{2^{N+2}}<t<$ 1, the Kepler problem $(M,\omega,F_{t}=(L,H_{t}))$ has at least $12^{N-1}+$ 1 nondisplaceable fibers.
@@
Now consider the symplectic manifold $(S^{2}\times S^{2},\omega_{S^{2}}\oplus\omega_{S^{2}})$ with Cartesian coordinates induced from ${\mathbb{R}}^{3}\times{\mathbb{R}}^{3}$, and the Hamiltonians $G:=(J,H):S^{2}\times S^{2}\rightarrow\mathbb{R}^{2}$, where
@@
Then $(S^{2}\times S^{2},\omega_{S^{2}}\oplus\omega_{S^{2}},G)$ is the Kepler problem for $R=1,t=1$ with a change of sign in the symplectic form. Set $S_{c}:=L^{-1}(0)\cap H^{-1}(c)=G^{-1}(0,c)$. Entov & Polterovich [entov2009rigid] constructed a partial symplectic quasi-state $\zeta_{-1}$ such that $S_{-1}$ is $\zeta_{-1}$-superheavy. Furthermore, according to Fukaya & Oh & Ohta & Ono [fukaya2011spectral], for every $c\in\ ]-1,-\frac{1}{2}]$, there exists a partial symplectic quasi-state $\zeta_{c}$ such that $S_{c}$ is $\zeta_{c}$-superheavy.
By Lemma 4.15, if $R\neq 0$, then all $\zeta_{c}$ with $c\in\left[-1,-\frac{1}{2}\right]$ are symplectic quasi-states for $(S^{2}\times S^{2},R(\omega_{S^{2}}\oplus\omega_{S^{2}}))$.
@@
By Theorem 2.31, there exists, for all $c\in A$, a pseudoheavy fiber $F_{t}^{-1}(p_{c})$ for some $p_{c}\in F_{t}(S^{2}\times S^{2})$ for the integrable systems $(S^{2}\times S^{2},R(\omega_{S^{2}}\otimes\omega_{S^{2}}),F_{t})$ with $0\leq t\leq 1$ and $R\neq 0$. In particular, by Theorem 2.30, a pseudoheavy fiber has to intersect the corresponding superheavy fiber, i.e., $F^{-1}_{t}(p_{c})\cap S_{c}\neq\emptyset$. Furthermore the set $F^{-1}_{t}(p_{c})$ is nondisplaceable. Now we have to ensure that the sets $F^{-1}_{t}(p_{c})$ are disjoint for different values of $c$. First notice that
@@
Hence, if $t>1-\frac{1}{2^{N+2}}$, for $c\in A$ the sets $[c-(1-t)4,\ c+(1-t)4]$ do not intersect each other. Therefore the nondisplaceable fibers $F_{t}^{-1}(p_{c})$ are disjoint. Hence we obtain the desired result. ∎
@@
The Lagrangian torus fibers under consideration are Chekanov-type tori, see [auroux2007mirror]. Furthermore due to the condition on $y$ the nondisplaceability follows from Vianna & Tonkonog [tonkonog2015low, Section $3$] or Auroux [auroux2007mirror], where the potential function of these tori is computed. For this choice of $y$ the potential function has a critical point, introducing a so called bulk deformation if necessary. For more details see for example Fukaya & Oh & Ohta & Ono [fukaya2012toric]. Therefore these tori are nondisplaceable. The focus-focus fiber is nondisplaceble since the above mentioned tori are nondisplaceable and since displaceability is an open property. ∎
@@
By Theorem 2.30 if the focus-focus fiber is $\zeta_{a}$-pseudoheavy then it intersects the superheavy fiber $S_{a}$. Therefore, there must exist a point $(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2})\in S^{2}\times S^{2}$ such that

### v2

@@ 3.1. Generalization of the method of probes
Let $F:(M,\omega)\rightarrow B$ be a toric fibration. Suppose that there exists an affine embedding of $([0,R]\times[0,h],\mathcal{A}_{0})$ into $(B,\mathcal{A})$ where $\mathcal{A}$ is the integral affine structure induced by $F$ on $B$, $R>0$ and $0<h<\frac{R}{2}$. Then the set $F^{-1}(\text{Im}([0,h]\times[0,h]))\subset M$ is displaceable.
@@
Let $F:(M,\omega)\rightarrow B$ be a toric fibration. Suppose that there exists an affine embedding of $([0,R]\times[0,a],\mathcal{A}_{0})$ into $(B,\mathcal{A})$ where $\mathcal{A}$ is the integral affine structure induced by $F$ on $B$, $R,a>0$ and $0<h<\frac{R}{2}$. Then the set $F^{-1}(\text{Im}([0,h]\times[0,a]))\subset M$ is displaceable.
@@ 3.2. Nondisplaceability of focus-focus fibers with higher multiplicity
Let $c$ be a focus-focus value of a semitoric system $F$ on a $4$-dimensional symplectic manifold $(M,\omega)$, and let $k$ be the number of focus-focus point on the fiber $F^{-1}(c)$, often called the multiplicity of the focus-focus fiber $F^{-1}(c)$. Intuitively, a fiber with $k$ focus-focus points is often seen as a torus with $k$ pinches. Alternatively the fiber $F^{-1}(c)$ can be seen as a closed chain of $k$ spheres joint at the poles, see Zung [Zun96] and Vũ Ngǫc [Vu00, Section $6$]. In particular, these are Lagrangian spheres. If $k=1$ the sphere is immersed, and if $k>1$ the spheres are embedded.
@@
Recall the notion of intersection number and Euler characteristic (for a reference see Hirsch [Hir94, Section $5.2$]).
Consider an embedded Lagrangian sphere $i:L\rightarrow(M,\omega)$ where $(M,\omega)$ is a $4$-dimensional symplectic manifold. Using Weinstein’s Lagrangian neighborhood theorem, we identify a neighborhood of $i(L)$ with the cotangent bundle of $\mathbb{S}^{2}$. Therefore the intersection number $L\cdot L$ is such that $|L\cdot L|=\chi(\mathbb{S}^{2})\neq 0$. Now we prove that $i(L)$ is not topologically displaceable by a smooth map $g:L\rightarrow M$ homotopic to the inclusion $i$. Let us argue by contradiction. Assume that $g(i(L))\cap i(L)=\emptyset$ and $g\simeq i$. In particular, $g$ and $i$ are transverse, and since $g(i(L))\cap i(L)=\emptyset$ we have $L\cdot L=\#(g,L)=0$, where #(g,L)\#(g,L) is the intersection number of the map $g$ with $L$. Therefore we obtain a contradiction. Since $L$ is not topologically displaceable, in particular it is not displaceable by a Hamiltonian diffeomorphism. ∎
@@
Therefore studying (non)displaceability of focus-focus fibers is reduced to the case $k=1$, which is the generic case, see Smirnov [Smi14]. There are also explicit examples for higher multiplicity. For the case $k=2$ see for instance De Meulenaere & Hohloch [DH21] and Section 4.3.
@@ 3.3. Displacing focus-focus fibers
Now assume that the corner introduced by the focus-focus fiber is a Delzant corner, and let $A$ be the integral affine transformation that takes it into the standard Delzant corner. By Remark 2.20 we can apply a nodal trade (which is defined in Lemma 2.15). Recall the affine distance from Equation (2.1) and denote the affine length by aff. Without loss of generality assume that $\text{aff}(v_{1})\geq\text{aff}(v_{2})$. Suppose that $h<\text{aff}(v_{2})$, and that there exists $R>0$ such that $R<\text{aff}(v_{1})$ and $0<2h<R$. Furthermore, suppose that the rectangle $0,R]\times[0,h]$ starting at $(0,0)$ sits inside $A(\Delta_{\vec{\epsilon}})$, as is sketched in Figure 3.2. Then we obtain the following result:
@@
Figure 3.2. Rectangle $0,R]\times[0,h]$ sitting inside $A(\Delta_{\vec{\epsilon}})$. The red dot represents the focus-focus value.
@@
In this section we consider (non)displaceability of fibers of the coupled spin-oscillator, Subsection 4.1, the coupled angular momenta, Subsection 4.2, and a semitoric system on the symplectic manifold given by the octagon, Subsection 4.3.
@@
Let $\rho_{1},\rho_{2}>0$ be positive constants. Consider the product manifold $M=\mathbb{S}^{2}\times\mathbb{R}^{2}$ with symplectic form $\omega=\rho_{1}\omega_{\mathbb{S}^{2}}\bigoplus\rho_{2}\omega_{\mathbb{R}^{2}}$, where $\omega_{\mathbb{S}^{2}}$ and $\omega_{\mathbb{R}^{2}}$ are the standard symplectic structures on the unit sphere and the Euclidean space respectively. Let $(x,y,z)$ be Cartesian coordinates on the unit sphere $S^{2}\subset\mathbb{R}^{3}$ and $(u,v)$ Cartesian coordinates on the plane $\mathbb{R}^{2}$.
@@
Coupled spin-oscillators are semitoric systems, for more details and background see Pelayo & Vũ Ngǫc [PV12]. The system has exactly one focus-focus singularity at the point $m:=(0,0,1,0,0)$. In order to understand the displaceability properties of the fibers of these systems we look at a representative of the polytope invariant, see Figure 4.1. These were computed in Pelayo & Vũ Ngǫc  [PV12] and Alonso [Alo19, Section $5.1$].
Figure 4.1. Representative of the polytope invariant of the coupled-spin-oscillator with $\vec{\epsilon}=+1$. The red dot represents the focus-focus value and is placed at $(0,0)$. Notice that the polytope is unbounded.
@@
Let $R_{1},R_{2}\in\mathbb{R}^{>0}$ and $t\in[0,1]$ a parameter. Consider the product manifold $M:=\mathbb{S}^{2}\times\mathbb{S}^{2}$ with symplectic form $\omega:=-(R_{1}\omega_{\mathbb{S}^{2}}\bigoplus R_{2}\omega_{\mathbb{S}^{2}})$ where $\omega_{\mathbb{S}^{2}}$ is the standard symplectic form on $\mathbb{S}^{2}$. Let $(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2})$ be Cartesian coordinates on $M$ induced from the ambient ${\mathbb{R}}^{3}\times{\mathbb{R}}^{3}$. The coupled angular momenta system is a family of $4$-dimensional completely integrable systems $(M,\omega,F_{t}:=(L,H_{t}))$, where the smooth functions $L,H_{t}:M\rightarrow\mathbb{R}$ are given by
@@
The system has four fixed points located at $(0,0,\pm 1,0,0,\pm 1)$. Three are of elliptic-elliptic type, namely $(0,0,\pm 1,0,0,1)$, but (0,0,1,0,0,1)(0,0,1,0,0,-1) is of focus-focus type for $t\in\ ]t^{-},t^{+}[$ and degenerate for $t\in\{t^{-},t{+}\}$, where
@@
Note that at the parameter $t=\frac{1}{2}$ there is always a focus-focus singularity. For more details and proofs of these results, see for instance Alonso [Alo19, Section 6] and Alonso & Dullin & Hohloch [ADH20].
@@
Note that the map $f$, defined as the composition of a horizontal translation with the straightening homeomorphism $f_{\vec{\epsilon}}$ of a representative of the polytope invariant, maps the vertical line $x=0$ in $F_{t}(M)$ to the vertical line $x=R_{2}-R_{1}$ in $f(F_{t}(M))$, see Alonso & Dullin & Hohloch [ADH20].
@@
Moreover, for $b\in\mathbb{R}^{>0}$, we have $b-2t,-b]\cap\{b\}=\emptyset$, hence the desired fibers are displaceable by $\Psi$. Note that the focus-focus value is $(0,1-2t)\in{\mathbb{R}}^{2}$. Hence, for $t<\frac{1}{2}$, the focus-focus fiber $F_{t}^{-1}(0,1-2t)$ is displaceable. ∎
Let us now inquire about the (non)displaceability of the focus-focus fiber for $t\geq\frac{1}{2}$. First we need to establish some facts about the height invariant. From now on $h:=h_{t}$ is the height invariant associated with the Kepler problem $(M,\omega,F_{t})$. In Alonso & Dullin & Hohloch [ADH20] the authors compute the height invariant of the Kepler problem and show that it is given by
@@
Figure 4.3. Plot of $\tilde{h}$ for $t\in\ ]t^{-},t^{+}[$ in blue. Plot of the constant line equal 11 in yellow. Plot of the constant line equal 0 in green.
@@
Step 1: Displaceability on the segment of the eigenline on the focus-focus value. Due to the choice of $t_{0}$ we have that for $t^{-}<t<t_{0}$, $\tilde{h}(t)>1$ and thus 2Rh<R2R-h<R. Hence we can apply Lemma 3.5 in Figure 4.5 to conclude that the focus-focus fiber and every fiber of the form $F_{t}^{-1}(0,b)$, where $(0,b)\in F_{t}(M)$ with b>12tb>1-2t, is displaceable.
@@
We are left with understanding the cases $t\geq t_{0}$. In Theorem 4.12 we will obtain a complete description of the (non)displaceability properties of the Kepler problem for $t=t_{0}$: there exists only one nondisplaceable fiber, and it is given by the focus-focus fiber, in particular it is a stem and hence nondisplaceable by symplectomorphisms. For $t>t_{0}$ we will see that there exist an infinite amount of nondisplaceable fibers, noticing that these fibers are equivalent to fibers already studied in the literature, see Auroux [Aur07]. In particular the focus-focus fiber is nondisplaceable. This is done in Section 4.2.3.
@@
In order to obtain further displaceability results we need to look at another representative of the polytope invariant for the Kepler problem. We use the one given by Figure 4.6, for more details see Alonso & Dullin & Hohloch [ADH20].
@@
Let $(0,y)\in f_{\vec{\epsilon}}(F_{t}(M))$ and $y>\frac{4}{3}R-h$. Notice that the vector $v:=(1,0)$ is integrally transverse to the facet given by y=x+(2Rh)y=x+(2R-h). Therefore we can construct the probe $(x+k,y)$ where y=x+(2Rh)>0y=x+(2R-h)>0. The affine distance of this probe is $2R-\frac{(y+h)}{2}$ since it will end at the facet given by y=2xhy=2x-h. Due to Lemma 2.22 the $(0,y)$ such that $(0,y)\in f_{\vec{\epsilon}}(F_{t}(M))$ and $y>\frac{4}{3}R-h$ give displaceable fibers, since they lie less than halfway along the probe. ∎
@@
Consider the Kepler problem for $t_{0}\leq t<1$. Let $f_{\vec{\epsilon}}$ be the straightening homeomorphism associated with the polytope invariant given by Figure 4.6. Then the fibers $F_{t}^{-1}(f_{\vec{\epsilon}}^{-1}(0,y))$ with $(0,y)\in f_{\vec{\epsilon}}(F_{t}(M))$ and y>Rhy>R-h of the Kepler problem are displaceable.
@@
The fibers we consider here correspond to fibers that sit above the segment of the eigenline $(x,x)$ in Figure 4.5 with x<2Rhx<2R-h. In the literature these are known as Chekanov-type tori, see Auroux [Aur07].
By applying a nodal slide, see Proposition 2.14, we can make the segment of the eigenline of Figure 4.6 as small as possible while preserving the fiber we are considering. This corresponds to considering the fiber $(x,x)$ in the polytope invariant, given by Figure 4.6, of the system for a $\tilde{t}>t$. Therefore the fiber over $(x,x)$ is displaceable by the probe that starts in the facet y=x+(2Rh)y=x+(2R-h) with direction (0,1)(0,-1). Hence we obtain our desired result. ∎
@@
Since the fiber for the Kepler problem $(M,\omega,F_{t_{0}})$ is actually monotone, one could use the methods of Hong & Kim & Lau [HKL23] or Rizzel & Ekholm & Tonkonog [RET22] to obtain the potential function for the focus-focus fiber, and then obtain a nondisplaceability result by computing the critical points of the potential function. However the result proven in Theorem 4.12 is stronger since we prove that the focus-focus fiber is a stem.
@@
Let $N\in\mathbb{N}$ and $R>0$. Then, for $1-\frac{1}{2^{N+2}}<t<1$, the Kepler problem $(M,\omega,F_{t}=(L,H_{t}))$ has at least $2^{N-1}+1$ nondisplaceable fibers.
@@
Now consider the symplectic manifold $(\mathbb{S}^{2}\times\mathbb{S}^{2},\omega_{\mathbb{S}^{2}}\oplus\omega_{\mathbb{S}^{2}})$ with Cartesian coordinates induced from ${\mathbb{R}}^{3}\times{\mathbb{R}}^{3}$, and the Hamiltonians $G:=(J,H):\mathbb{S}^{2}\times\mathbb{S}^{2}\rightarrow\mathbb{R}^{2}$, where
@@
Then $(\mathbb{S}^{2}\times\mathbb{S}^{2},\omega_{\mathbb{S}^{2}}\oplus\omega_{\mathbb{S}^{2}},G)$ is the Kepler problem for $R=1,t=1$ with a change of sign in the symplectic form. Set $S_{c}:=L^{-1}(0)\cap H^{-1}(c)=G^{-1}(0,c)$. Entov & Polterovich [EP09] constructed a partial symplectic quasi-state $\zeta_{-1}$ such that $S_{-1}$ is $\zeta_{-1}$-superheavy. Furthermore, according to Fukaya & Oh & Ohta & Ono [FOOO11b], for every $c\in\ ]-1,-\frac{1}{2}]$, there exists a partial symplectic quasi-state $\zeta_{c}$ such that $S_{c}$ is $\zeta_{c}$-superheavy.
By Lemma 4.15, if $R\neq 0$, then all $\zeta_{c}$ with $c\in\left[-1,-\frac{1}{2}\right]$ are partial symplectic quasi-states for $(\mathbb{S}^{2}\times\mathbb{S}^{2},R(\omega_{\mathbb{S}^{2}}\oplus\omega_{\mathbb{S}^{2}}))$.
@@
By Theorem 2.31, there exists, for all $c\in A$, a pseudoheavy fiber $F_{t}^{-1}(p_{c})$ for some $p_{c}\in F_{t}(\mathbb{S}^{2}\times\mathbb{S}^{2})$ for the integrable systems $(\mathbb{S}^{2}\times\mathbb{S}^{2},R(\omega_{\mathbb{S}^{2}}\otimes\omega_{\mathbb{S}^{2}}),F_{t})$ with $0\leq t\leq 1$ and $R\neq 0$. In particular, by Theorem 2.30, a pseudoheavy fiber has to intersect the corresponding superheavy fiber, i.e., $F^{-1}_{t}(p_{c})\cap S_{c}\neq\emptyset$. Furthermore the set $F^{-1}_{t}(p_{c})$ is nondisplaceable. Now we have to ensure that the sets $F^{-1}_{t}(p_{c})$ are disjoint for different values of $c$. First notice that
@@
Hence, if $t>1-\frac{1}{2^{N+2}}$, for $c\in A$ the sets [c(1t)4,c+(1t)4][c-(1-t)4,\ c+(1-t)4] do not intersect each other. Therefore the nondisplaceable fibers $F_{t}^{-1}(p_{c})$ are disjoint. Hence we obtain the desired result. ∎
@@
The Lagrangian torus fibers under consideration are Chekanov-type tori, see [Aur07]. Furthermore due to the condition on $y$ the nondisplaceability follows from Vianna & Tonkonog [TV18, Section $3$] or Auroux [Aur07], where the potential function of these tori is computed. For this choice of $y$ the potential function has a critical point, introducing a so called bulk deformation if necessary. For more details see for example Fukaya & Oh & Ohta & Ono [FOOO12]. Therefore these tori are nondisplaceable. The focus-focus fiber is nondisplaceble since the above mentioned tori are nondisplaceable and since displaceability is an open property. ∎
@@
By Theorem 2.30 if the focus-focus fiber is $\zeta_{a}$-pseudoheavy then it intersects the superheavy fiber $S_{a}$. Therefore, there must exist a point $(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2})\in\mathbb{S}^{2}\times\mathbb{S}^{2}$ such that